# Demonstration of weighted graph optimization on a Rydberg atom array using local light-shifts

### v2

@@ Abstract
Discussion and Outlook.— In this paper we have demonstrated a new approach to implementing UDG-MWIS using arbitrary local light-shifts implemented using a SLM combined with a dual stage annealing profile that combines the advantages of local light-shifts with global frequency and intensity control for implementing a quantum annealing algorithm. We demonstrate the ability to prepare ground states of small-scale 1D and 2D graphs, including demonstrating the embedding of a non-UDG MWIS problem. In both cases, we were able to identify robust annealing protocols suitable for use with different relative weightings providing the possibility to design universal ramps that are suitable for solving a large number of graphs without the required overhead of full closed-loop optimization of the annealing ramp each time.
@@
This approach for local light-shifts is highly scalable, with SLMs able to generate arrays of $>\!\!1000$ individually controllable spots and the option to implement in-situ realtime feedback using a camera calibrated against the measured atomic shift previously demonstrated for normalization of arbitrary trapping potentials [25]. Currently at 800 nm we have up to 4 W available at the experiment, providing sufficient power for $>\!\!100$ weighted sites. In future, this can be scaled further using coherent combination or frequency doubling telecoms fiber lasers to increase the available power and combined with techniques for beam-shaping to implement a flat-top point spread function on the SLM [57] to reduce the size of the light-shift spots and further reduce mechanical effects caused by using a repulsive optical potential to enable longer annealing runs.
@@ Acknowledgements.
Supplemental Material for Demonstration of weighted graph optimization on a Rydberg atom array using local light-shifts
@@
I Experiment Setup
I.1 Neutral Atom Arrays
A schematic of the experimental setup is shown in Fig. 6(a), based on the setup previously introduced in Ref. [20]. Atom arrays are created by loading individual Cs atoms into 1064 nm optical tweezers generated using a spatial light modulator (SLM). Traps are focused to a $1/e^{2}$ waist of 1.5 $\mu$m using an in-vacuum aspheric lens with NA=0.45 [60] which has an ITO coating to suppress charge build up. This same lens is used for qubit readout by performing fluorescence imaging with an sCMOS (Teledyne Photometrics Prime BSI) camera.
@@ I.1 Neutral Atom Arrays
For the experiments in this paper, prior to Rydberg operations the optical dipole traps are ramped down to 15 $\mu$K, cooling the atoms to temperatures of 2 $\mu$K. Atoms are then released for 5 $\mu$s to minimise the mechanical force of the tweezer beams on the Rydberg states, and atoms are then recaptured at 1 mK to facilitate efficient Rydberg state detection [61].
I.2 Rydberg Excitation
Rydberg excitation is performed using two-photon excitation via the intermediate excited state as shown in Fig. 6(b) using light at 459 nm and 1039 nm to drive transitions from $6P_{1/2}\rightarrow 7P_{1/2}$ and $7P_{1/2}\rightarrow nS_{1/2}$ respectively. Both beams are derived from Ti:Sapph lasers (M Squared SolsTiS) at 918 nm and 1039 nm which are frequency stabilised to a commercial ULE cavity (M Squared) offering finesse $\mathcal{F}>40$ k at both wavelengths. The lasers are offset-locked using a Pound-Drever-Hall with fiber EOMs [62] to provide controllable detuning with respect to the 3 GHz spaced cavity modes, with few kHz linewidths measured via a beatnote between the two lasers.
The 459 nm light is frequency doubled (M Squared ECDF) and the detuning controlled using a double pass accousto-optic modulator (AOM,AA MQ180-A0.25-VIS) in a cat-eye geometry to enable fast frequency control. The AOM is driven using a high bandwidth arbitrary waveform generator (AWG, Spectrum M4i6631-x8). This is delivered to the experiment by fiber, and overlapped onto the array with a 125 $\mu$m $1/e^{2}$ waist. The 1039 nm laser seeds a fiber amplifier (Azur ALS-1039-20-CP) which is intensity controlled using a single-pass AOM and coupled into a high-power fiber. The beam is then imaged onto the atoms using cylindrical optics to create a beam with a $1/e^{2}$ waist size of $60\times 20~{}\mu$m. For the experiments presented here, Rydberg excitation is performed using an intermediate detuning of $\Delta^{\prime}/2\pi=506$ MHz with respect to the $7P_{1/2}$ centre of mass energy, with powers set to provide a Rabi frequency of $\Omega/2\pi=2.70(0)$ MHz over the central rows of the array.
@@ I.2 Rydberg Excitation
I.3 Local Light-shifts
To apply local light-shifts for implementing graph weighting, an 800 nm laser (M Squared SolsTiS) is used with a second spatial light modulator to create a secondary tweezer array that is overlapped with the underlying 1064 nm traps as shown in Fig. 6(a). Due to the chromatic shifts of the in-vacuum aspheric, it is not possible to image both 1064 nm and 800 nm arrays overlapped after the chamber focused in a common plane. Instead, to overlap the arrays we exploit the fact that ground-state atoms are anti-trapped by this blue detuned light. Using this effect, we initially generate arrays with a 2 $\mu$m waist and apply the 800 nm potential onto the trapped atoms for 1-10 ms, resulting in ejection of atoms from sites that are overlapped between the two beams. Using $10\times 10$ arrays with 8 $\mu$m spacing, we perform multiple iterations to adjust relative spacing, rotation and defocus of the 800 spots with respect to the 1064 nm array using the SLM to maximise the relative alignment of the two arrays by minimising atom survival. This also ensures the 800 nm light is focused on the same plane as the 1064 nm traps.
For performing local-light shifts, we suppress the mechanical effects of the repulsive light shift on the ground-state atoms by changing the input beam size on the SLM to create spots with an effective waist size of 3.0(4) $\mu$m, resulting in $<2\%$ additional atom loss when applying local shifts of up to 10 MHz for periods of up to 3 $\mu$s.
@@ I.3 Local Light-shifts
I.4 Weighted-graph Annealing
@@
To synchronise the pulses, the parameters are sent to the same dual channel AWG with one channel controlling amplitude and frequency of the 459 nm laser, and the second channel connected to the 800 nm AOM situated before the SLM that provides intensity control.

### v3

@@ V.3 Crossing Gadget
As discussed above, a number of proposals exist to extend the range of target optimisation problems that can be solved using neutral atom arrays by converting to a UDG-MWIS [45, 46, 47]. One such approach introduces simple gadgets that provide a generalised framework for mapping problems to UDG-MWIS [46], and here we consider a proof-of-concept implementation of the crossing gadget using our local light-shift approach. This crossing gadget, shown in Fig. 5(a), implements an edge between qubits 1 and 2, and qubits 3 and 4. This requires four additional atoms, where the logical qubits (1-4) are assigned weight $w_{1}=1$, whilst the ancilla gadgets have $w_{2}=2.86$ to ensure the lowest energy state corresponds to an equal superposition of the four degenerate output configurations with either qubit 1 or 2 excited independent of 3 or 4. This deviates from $w_{2}=4$ presented in [46] due to the requirement to compensate for the effects of the $1/R^{6}$ interaction tails on the ancilla atoms (See App. C). For these measurements the intermediate state detuning was increased to $\Delta^{\prime}/2\pi=1$ GHz and Rabi frequency reduced to $\Omega/2\pi=1.28$ MHz motivated by numerical modelling of the annealing process. As above, we then optimize the annealing ramp experimentally to minimize $\langle H_{\mathrm{MWIS}}\rangle$ for the problem resulting in the ramp shown in Fig. 5(b). The distribution of the final output strings are shown in Fig. 5(c), where the four target ground states are the most likely to be observed with approximately equal probability, and $\geq 20\%$ combined probability. Analysis of the next four most probable output strings show these correspond to the case of a target ground state with one of the logical qubit Rydberg excitation’s missing, likely due to limitations in the Rydberg detection fidelity we estimate a detection error of 8.5% for the current detection method based on ejection from the tweezer potentials [63]). For embedding of larger graphs, improvements in the Rydberg state detection can be implemented using microwave ionization [64, 29], however these results verify the proof-of-concept approach for performing gadget encodings on the current architecture.
VI Discussion and Outlook
@@
In this paper we have demonstrated a new approach to implementing UDG-MWIS using arbitrary local light-shifts implemented using a SLM combined with a dual stage annealing profile that combines the advantages of local light-shifts with global frequency and intensity control for implementing a quantum annealing algorithm. We demonstrate the ability to encode ground states of small-scale 1D and 2D graphs, including demonstrating the embedding of a non-UDG MWIS problem. In both cases, we were able to identify robust annealing protocols suitable for use with different relative weightings providing the possibility to design universal ramps that are suitable for solving a large number of graphs without the required overhead of full closed-loop optimization of the annealing ramp each time.
@@
This approach for local light-shifts is highly scalable, with SLMs able to generate arrays of $>\!\!1000$ individually controllable spots and the option to implement in-situ realtime feedback using a camera calibrated against the measured atomic shift previously demonstrated for normalization of arbitrary trapping potentials [25]. Currently at 800 nm we have up to 4 W available at the experiment, providing sufficient power for $>\!\!100$ weighted sites. In future, this can be scaled further using coherent combination or frequency doubling telecoms fiber lasers to increase the available power and combined with techniques for beam-shaping to implement a flat-top point spread function on the SLM [65] to reduce the size of the light-shift spots and further reduce mechanical effects caused by using a repulsive optical potential to enable longer annealing runs. For larger graphs the system also requires increased uniformity of Rydberg excitation over the array which can be implemented using additional SLMs in the global Rydberg beam paths to create top-hat intensity profiles [29].
The measured output state probabilities in the current experiments are limited to 10-20% by a number of factors. Fundamental limits arise from the finite-duration of the annealing profile and spontaneous decay from the Rydberg state alongside scattering from the intermediate excited state, with numerical simulations predicting target output probabilities between 50-60% for the graphs used here. This is further reduced by the effects of finite state preparation fidelity, finite Rydberg detection fidelity and additional losses such as finite vacuum lifetime. By implementing enhanced optical pumping techniques [21], microwave-assisted Rydberg detection [64, 29], extended vacuum lifeitmes [19] and increased laser power to suppress losses by facilitating larger intermediate state detunings [22] these limitations can be suppressed to maintain high-fidelity readout for large arrays.
@@ Acknowledgements.
Appendix A Experiment Setup
Neutral Atom Arrays. — A schematic of the experimental setup is shown in Fig. 6(a), based on the setup previously introduced in Ref. [20]. Atom arrays are created by loading individual Cs atoms into 1064 nm optical tweezers generated using a spatial light modulator (SLM). Traps are focused to a $1/e^{2}$ waist of 1.5 $\mu$m using an in-vacuum aspheric lens with NA=0.45 [68] which has an ITO coating to suppress charge build up. This same lens is used for qubit readout by performing fluorescence imaging with an sCMOS (Teledyne Photometrics Prime BSI) camera.
@@ Appendix A Experiment Setup
For the experiments in this paper, prior to Rydberg operations the optical dipole traps are ramped down to 15 $\mu$K, cooling the atoms to temperatures of 2 $\mu$K. Atoms are then released from the traps for 5 $\mu$s to minimise the mechanical force of the tweezer beams on the Rydberg states, and atoms are then recaptured at 1 mK to facilitate Rydberg state detection via ejection of the Rydberg states from the traps. For the current parameters we estimate a Rydberg detection error of 8.5% [63].
Rydberg Excitation. — Rydberg excitation is performed using two-photon excitation via the intermediate excited state as shown in Fig. 6(b) using light at 459 nm and 1039 nm to drive transitions from $6P_{1/2}\rightarrow 7P_{1/2}$ and $7P_{1/2}\rightarrow nS_{1/2}$ respectively. Both beams are derived from Ti:Sapph lasers (M Squared SolsTiS) at 918 nm and 1039 nm which are frequency stabilised to a commercial ULE cavity (M Squared) offering finesse $\mathcal{F}>40$ k at both wavelengths. The lasers are offset-locked using a Pound-Drever-Hall with fiber EOMs [69] to provide controllable detuning with respect to the 3 GHz spaced cavity modes, with few kHz linewidths measured via a beatnote between the two lasers.
The 459 nm light is frequency doubled (M Squared ECDF) and the detuning controlled using a double pass accousto-optic modulator (AOM,AA MQ180-A0.25-VIS) in a cat-eye geometry to enable fast frequency control. The AOM is driven using a high bandwidth arbitrary waveform generator (AWG, Spectrum M4i6631-x8). This is delivered to the experiment by fiber, and overlapped onto the array with a 125 $\mu$m $1/e^{2}$ waist. The 1039 nm laser seeds a fiber amplifier (Azur ALS-1039-20-CP) which is intensity controlled using a single-pass AOM and coupled into a high-power fiber. The beam is then imaged onto the atoms using cylindrical optics to create a beam with a $1/e^{2}$ waist size of $60\times 20~{}\mu$m. For the experiments presented here, Rydberg excitation is performed using an intermediate detuning of $\Delta^{\prime}/2\pi=502$ MHz with respect to the $7P_{1/2}$ centre of mass energy, with powers set to provide a Rabi frequency of $\Omega/2\pi=2.70(0)$ MHz over the central rows of the array.
@@
Local Light-shifts. — To apply local light-shifts for implementing graph weighting, an 800 nm laser (M Squared SolsTiS) is used with a second spatial light modulator to create a secondary tweezer array that is overlapped with the underlying 1064 nm traps as shown in Fig. 6(a). Due to the chromatic shifts of the in-vacuum aspheric, it is not possible to image both 1064 nm and 800 nm arrays overlapped after the chamber focused in a common plane. Instead, to overlap the arrays we exploit the fact that ground-state atoms are anti-trapped by this blue detuned light. Using this effect, we initially generate arrays with a 2 $\mu$m waist and apply the 800 nm potential onto the trapped atoms for 1-10 ms, resulting in ejection of atoms from sites that are overlapped between the two beams. Using $10\times 10$ arrays with 8 $\mu$m spacing, we perform multiple iterations to adjust relative spacing, rotation and defocus of the 800 spots with respect to the 1064 nm array using the SLM to maximise the relative alignment of the two arrays by minimising atom survival. This also ensures the 800 nm light is focused on the same plane as the 1064 nm traps.
For performing local-light shifts, we suppress the mechanical effects of the repulsive light shift on the ground-state atoms by changing the input beam size on the SLM to create spots with an effective waist size of 3.0(4) $\mu$m, resulting in $<2\%$ additional atom loss when applying local shifts of up to 10 MHz for periods of up to 3 $\mu$s. In the present work this is achieved by masking the active area of the SLM to reduce the effective beam size at the cost of reduced efficiency and limitations in the number of spots that can be generated due to finite SLM pixels. In future this limitation can be readily overcome by adapting the magnification of the relay optics from SLM to the atoms enable high-efficiency generation of arrays of over 1000 spots [61].
@@
Note that this same SLM based approach could be used to create local light shifts using red-detuned light, with the requirement that now vertices with low weight require higher power light shifts and global detuning and light-shift beams need to be modulated simultaneously during the positive detuning part of the ramp. Compared to our implementation with blue-detuned light, this approach both places greater demands on the bandwidth of the AOM driving the global detuning and also couples the effects of a finite AOM bandwidth and finite local light-shift beam power in determining the maximum achievable total detuning. The use of red-detuned light shift beams is thus more technically challenging and, while mitigating the anti-trapping effect of the light shift beams on the atoms in their ground state, still has an anti-trapping effect on the Rydberg state atoms
Appendix B Weighted-graph Annealing
@@
To synchronise the pulses, the parameters are sent to the same dual channel AWG with one channel controlling amplitude and frequency of the 459 nm laser, and the second channel connected to the 800 nm AOM situated before the SLM that provides fast intensity control.
@@
Figure 8: Crossing gadget weightings. (a) Crossing gadget schematic showing labels for weight and qubit number. (b) One of the four possible target output states that encodes the UDG solution (c) Actual ground-state for the case of uniform spacings (i.e. $r_{15}=r_{56}$), $w_{1}=1$ and $w_{2}=4$ when including real $1/R^{6}$ interaction tails.
Appendix C Crossing Gadget Weightings
In Sec. V.3 above we demonstrate use of the crossing gadget initially introduced by [46] and shown schematically in Fig. 8(a) with qubit labels. In the UDG limit, where we consider only the interactions between qubits connected by edges, the target ground-state can be encoded using weights $w_{1}=1$ and $w_{2}=4$ resulting in a cost of $C=-2w_{1}-w_{2}=-6$ for each of the four degenerate output states as demonstrated in Fig. 8(b).
When embedded on the realistic neutral atom system however with interactions of the form $V(R)=-C_{6}/R^{6}$, the energy of this target configuration becomes $E_{0}=-2w_{1}\delta_{0}-w_{2}\delta_{0}+V(r_{13})+V(r_{17})+V(r_{37})$. In the case that all edges are equal distance, $w_{1}=1$ and $w_{2}=4$, then the true groundstate of the system becomes that shown in Fig. 8(c) with energy $E^{\prime}=-2w_{2}\delta_{0}+V(r_{57})$.
To mitigate this issue and ensure $E_{0}<E^{\prime}$, we adjust the spacings to make $r_{75}\simeq r_{15}$ and choose $w_{2}$ to maximise the energy gap between $E_{0}$ resulting in choice of $w_{2}=2.86$ and $r_{75}=0.8r_{15}$ in the experiments above.